NUCLEAR SPIN MODIFICATION OF METHANE IN PARAHYDROGEN CRYSTALS

\underline{MIZUHO FUSHITANI}, MASAAKI MIKI, and TAKAMASA MOMOSE, Division of Chemistry, Graduate School of Science, Kyoto University, Kyoto 606-8502, JAPAN.

Solid parahydrogen is an excellent matrix for matrix-isolation spectroscopy because of its high spectral resolution. Here, we report the nuclear spin symmetry selection rule of a chemical reaction CH_{2} + H_{2} → CH_{4} and the nuclear spin conversion of CH_{4} embedded in parahydrogen crystals studied by infrared absorption spectroscopy. Due to the nuclear spin modification, not only the J=0 rotational level but also the J=1 level of CH_{4} is populated even if the temperature is lowered sufficiently. Since conversion among different nuclear spin states is very slow, population ratio between the J=1 and J=0 levels observed just after the crystal growth reflects the ratio of A and F nuclear spin states before the cooling. The vibration-rotation absorption of CH_{4} in parahydrogen crystal grown from a premixed gas of methane and hydrogen molecules shows the population ratio of approximately (J=1):(J=0)=2:1. On the other hand, CH_{4} produced by the photochemical reaction of CH_{3} + H_{2} + 193nm → CH_{2} + H + H_{2} → CH_{4} + H was found to have four times larger population in the J=1 rotational level than in the J=0 level just after the reaction. The former reflects the ratio of the A and F nuclear spin states being 5:9 at room temperature, while the latter is due to a nuclear spin selection rule of the reaction. In addition, the vibration-rotation absorptions of CH_{4} exhibit time-dependent intensity changes at 4.8K. These changes are interpreted to be a result of the I = 1(F) → I = 2(A) nuclear spin conversion which accompanies the J = 1 → J = 0 rotational relaxation. The half-lifetime of the upper J = 1 rotational state is unchanged by the addition of up to 2% orthohydrogen molecules, but decreases with more than 10% orthohydrogen molecules. The increase of the decay rate at higher orthohydrogen concentration indicates that the magnetic field gradient across CH_{4} caused by orthohydrogen molecules mixes the nuclear spin states which accelerate the conversion.